OpenStudy (anonymous):

can someone please help Factor the following expression 25w^2+30w+9

OpenStudy (shamim):

25w^2+30w+9 =25w^2+15w+15w+9 =5w(5w+3)+3(5w+3) =(5w+3)(5w+3)

OpenStudy (anonymous):

ty